# Revolutionizing Undergraduate Learning: CourseGPT and Its Generative AI Advancements

### v1

@@ I Introduction
This work evaluates RAG-based-LLM effectiveness as intelligent university course assistants. We aim to assess how different LLM sizes, ranging from 7 billion to 47 billion parameters, impact their ability to assist students with questions regarding the course. We identify the strengths and limitations of each LLM by establishing a shared knowledge base and employing rigorous evaluation methodologies. LLM assessment includes analyzing the correctness, context recall, and faithfulness of responses generated by each LLM within the course’s scope. CourseGPT utilizes Mistral LLMs [1, 2].
@@ II-A CourseGPT Impact
II-B Enhancing of Students’ Learning Outcomes
CourseGPT will serve as a valuable tool for instructors and students, providing specific guided procedures to implement three student outcomes from ABET: SO1, SO4, and SO7, thereby enhancing specific learning outcomes. This will guide us in designing the tool while offering instructors the necessary procedures to implement them, resulting in an improved student learning experience, regardless of whether the class is intended for ABET accreditation.
@@ II-B1 SO1: An ability to identify, formulate, and solve complex engineering problems
CourseGPT assists instructors in breaking down their material into three phases. Students will engage in broad, targeted searches, asking multiple questions to identify problems. They will then receive feedback on formulating these problems. Solving the problems requires students to ask more questions and seek similar solutions to validate their own. These activities benefit from CourseGPT’s interactive and individualized feedback during real-time sessions.
@@ II-B2 SO4: An ability to recognize ethical and professional responsibilities
Drawing from years of teaching experience, this outcome greatly benefits from CourseGPT by encouraging critical thinking. Linking a project’s impact to global, economic, environmental, and societal contexts requires long discussions and repeated, focused questions. CourseGPT’s real-time, interactive sessions of intelligent discussions significantly enhance students’ understanding of their project’s impact on its environment.
@@ II-B3 SO7: An ability to acquire and apply new knowledge
Achieving this ability goes beyond class material. Traditional instruction methods are limited to classrooms and the overwhelming information on the internet. CourseGPT allows instructors to compile supplemental material, facilitating interactive discussions between students and CourseGPT. This precise guidance helps students acquire and apply new knowledge effectively.
In summary, CourseGPT represents a significant advancement in educational technology, providing personalized, real-time support that enhances undergraduate learning experiences and outcomes. By facilitating complex problem-solving, promoting ethical responsibilities, and encouraging new knowledge acquisition, CourseGPT aligns with every school’s strategic educational goals. The pilot phase will establish a foundation for broader implementation, benefiting a larger cohort of students and setting a new standard for educational excellence in higher education.
@@ V-D Survey Results
The survey assesses the satisfaction and performance of CourseGPT based on the feedback from six participants comprising former students and an experienced teaching assistant. The survey result analysis unveils valuable insights into CourseGPT’s perceived effectiveness and utility in educational contexts. The survey questions were scored based on the 5-point Likert Scale.
@@ V-E Result Discussions
Analyzing different-sized LLMs offers valuable insights for optimizing workflows and improving retrieval support mechanisms for students enrolled in a course. The analysis highlights significant variances in correctness, context recall, and faithfulness scores among the LLMs, all determining CourseGPT’s effectiveness in research tasks and facilitating efficient experimentation.
The larger model, Mixtral-8x7b, which achieved higher correctness, context, and faithfulness scores, exemplifies superior proficiency in delivering precise and contextually relevant information. This precision is critical for experimenters, enhancing their efficiency by enabling quicker access to necessary information and supporting informed decision-making during the setup and execution of experiments.
These advantages enhance workflow efficiency and improve retrieval aids for students. By utilizing LLMs optimized for high correctness, context recall, and faithfulness to the source material, CourseGPT helps streamline workflow processes, expedite experiment setup and configuration, and reduce the likelihood of errors or inaccuracies. CourseGPT is a gateway for quickly accessing and retrieving course-related information, allowing students to gain more course-related quicker. Table 7 shows an example output from the highest-overall performer, Mixtral8x7b.
@@ VI-C Data Quality and Relevance
One of the key obstacles we face is guaranteeing the quality and appropriateness of the data CourseGPT employs for knowledge retrieval and creation. The precision and scope of the class documentation, from which the data is directly extracted, profoundly affects CourseGPT’s ability to offer valuable aid to students.

### v2

@@ I Introduction
LLMs are known to have broad knowledge on various corpus of data, unless they have been fine-tuned. Fine-tuning LLMS is an exhaustive and time-consuming process. A method substituting fine-tuning methods is Retrieval Augmented Generation (RAG) which provides a knowledge base for the LLM to utilize in answering prompts [1]. This work evaluates RAG-based-LLM effectiveness as intelligent university course assistants. We aim to assess how different LLM sizes, ranging from 7 billion to 47 billion parameters, impact their ability to assist students with questions regarding the course. We identify the strengths and limitations of each LLM by establishing a shared knowledge base and employing rigorous evaluation methodologies. LLM assessment includes analyzing the correctness, context recall, and faithfulness of responses generated by each LLM within the course’s scope. CourseGPT utilizes Mistral LLMs [2, 3].
@@ II-A CourseGPT Impact
II-B Enhancing Students’ Learning Outcomes
CourseGPT is designed to be a valuable tool for both instructors and students by providing guided procedures to implement three specific student outcomes from ABET: SO1, SO4, and SO7. This tool aims to enhance learning outcomes and also offers structured guidance for instructors, improving the overall student learning experience, whether or not the class seeks ABET accreditation.
@@ II-B1 SO1: An Ability to Identify, Formulate, and Solve Complex Engineering Problems
CourseGPT assists instructors in structuring their material into three phases: identification, formulation, and solution of problems. Initially, students engage in broad, targeted searches and ask multiple questions to identify problems. Next, they receive detailed feedback to help formulate these problems accurately. Finally, students are guided to solve these problems by asking further questions and exploring similar solutions to validate their approaches. CourseGPT’s interactive, real-time feedback and individualized support during these sessions enhance students’ problem-solving skills and deepen their understanding of complex engineering challenges.
@@ II-B2 SO4: An Ability to Recognize Ethical and Professional Responsibilities
Leveraging years of teaching experience, CourseGPT significantly enhances students’ ability to recognize ethical and professional responsibilities. It promotes critical thinking by linking a project’s impact to global, economic, environmental, and societal contexts. Through long discussions and repeated, focused questions, CourseGPT facilitates intelligent, real-time interactive sessions that deepen students’ understanding of the broader implications of their projects. This structured approach helps students develop a more comprehensive perspective on their ethical and professional responsibilities.
@@ II-B3 SO7: An Ability to Acquire and Apply New Knowledge
Achieving this outcome requires going beyond traditional classroom instruction. CourseGPT addresses this need by allowing instructors to compile supplemental material and facilitate interactive discussions between students and the tool. This capability helps students navigate the vast amount of information available on the internet and apply new knowledge effectively. CourseGPT’s precise guidance and real-time support enable students to acquire and integrate new knowledge seamlessly into their existing understanding.
In summary, CourseGPT represents a significant advancement in educational technology, providing personalized, real-time support that enhances undergraduate learning experiences and outcomes. By facilitating complex problem-solving, promoting ethical responsibilities, and encouraging the acquisition of new knowledge, CourseGPT aligns with strategic educational goals. The pilot phase will establish a foundation for broader implementation, benefiting a larger cohort of students and setting a new standard for educational excellence in higher education.
@@ V-D Survey Results
The survey assesses the satisfaction and performance of CourseGPT based on the feedback from six participants comprising former students and an experienced teaching assistant. The survey result analysis unveils valuable insights into CourseGPT’s perceived effectiveness and utility in educational contexts. The survey questions were scored based on the 5-point Likert Scale. The survey consisted of 12 questions and were then combined into categories, as seen in the results in Figs. 6, 6, 6, for simplicity.
@@ V-E Result Discussions
Analyzing different-sized LLMs offers valuable insights for optimizing workflows and improving retrieval support mechanisms for students enrolled in a course. The analysis highlights significant variances in correctness, context recall, and faithfulness scores among the LLMs, all determining CourseGPT’s effectiveness in answering course-related queries.
The larger model, Mixtral-8x7b, which achieved higher correctness, context, and faithfulness scores, exemplifies superior proficiency in delivering precise and contextually relevant information. This precision is critical for students, enhancing their efficiency by enabling quicker access to necessary information and supporting informed decisions regarding their study materials and habits.
These advantages enhance workflow efficiency and improve retrieval aids for students. By utilizing LLMs optimized for high correctness, context recall, and faithfulness to the source material, CourseGPT helps streamline workflow processes, and reduce the likelihood of errors or inaccuracies in answers. CourseGPT is a gateway for quickly accessing and retrieving course-related information, allowing students to gain more course-related information quicker. Fig. 7 shows an example output from the highest-overall performer, Mixtral8x7b.
@@
The implementation and testing of CourseGPT have yielded several significant implications for enhancing student learning outcomes and educational practices:
V-E1 Enhanced Instructor Efficiency
Instructors benefit from the structured guidance CourseGPT offers, allowing them to focus more on high-level pedagogical strategies rather than routine question answering. This leads to more efficient use of class time and better allocation of instructional resources, ultimately improving the overall quality of education.
V-E2 Personalized Learning Experiences
CourseGPT’s ability to provide individualized feedback and tailored learning pathways supports diverse student needs, leading to more personalized learning experiences. This adaptability helps in addressing different learning paces and styles, contributing to more inclusive and effective education.
The results demonstrate that CourseGPT significantly enhances student engagement, performance, and personalized learning experiences while improving instructor efficiency and promoting critical thinking. These implications highlight the potential of CourseGPT to transform educational practices and contribute to higher standards of education.
@@ VI-C Data Quality and Relevance
A key obstacle we face is guaranteeing the quality and appropriateness of the data CourseGPT employs for knowledge retrieval and creation. The precision and scope of the class material, from which the data is directly extracted, directly affects CourseGPT’s ability to offer valuable aid to students.